An lpg cylinder containing 15 kg butane at 27 degree Celsius and 10 atm pressure is leaking.After 1 day the pressure is decreased to 8 atm the quantity of gas leaked is

Grade:11

2 Answers

Vikas TU
7 years ago
This would be solved by using simple ideal gas equation that is PV = nRT.
PV = mRT
P1/m1 = P2/m2
Now putting the values we get,
10/15 = 8/m2
we get, m2 = 13 gm.
the gas leaked would be obvious =>  15 – 13 => 2 gm.
KAVINILAVU
4 years ago
pv=nrt
p1/m1=p2/m2
Now by substituing the values we get,
10/15=8/m2
we get, m2=12
the gas leaked would be => 15-12=3